OpenStudy (anonymous):

Help me please! Will fan and medal! 2x-3y=-6 y=-3x+2

TheSmartOne (thesmartone):

Which method would you want to use to solve this? Substitution Eliminator Graphing ? (:

OpenStudy (anonymous):

Substitution

TheSmartOne (thesmartone):

That's a good choice! :)

OpenStudy (anonymous):

Thanks

TheSmartOne (thesmartone):

2x-3y=-6 y=-3x+2 So we take the second equation y = -3x + 2 and we substitute it in to 2x - 3y = -6

TheSmartOne (thesmartone):

Do you know how we would substitute the 2nd formula into the 1st one, or do you need me to show you? :)

OpenStudy (anonymous):

I think so do you use the distributive property?

TheSmartOne (thesmartone):

Kind of. :)

TheSmartOne (thesmartone):

2x - 3y = -6 so we have a y value in our first equation the second equation is y = -3x + 2 we just switch the y in the first equation with -3x + 2

OpenStudy (anonymous):

@TheSmartOne so what is the answer?

TheSmartOne (thesmartone):

2x - 3(-3x + 2) = -6 Solve for x

OpenStudy (anonymous):

@TheSmartOne the final answer is (0,2)

TheSmartOne (thesmartone):

correct